# DAE-Fuse: An Adaptive Discriminative Autoencoder for Multi-Modality Image Fusion

### v1

@@ III-B1 Qualitative comparisons
As shown in Figure 2, the selected scenario highlights feature extraction and model bias. DDcGAN and TarDAL, both GAN-based models, exhibit noticeable blurriness and detail loss. DIDFuse, an AE-based method, shows a clear bias towards infrared images, darkening the sky and grass. Our DAE-Fuse, however, delivers the best results, preserving rich textural details and balancing both input modalities seamlessly. Notably, it maintains intact roof textures, unlike other methods. CDDFuse, which also uses a parallel encoder, partially retains roof details but overexposes some areas and introduces extra noise on the house wall. In contrast, DAE-Fuse fuses wall textures naturally from both inputs.
III-B2 Quantitative comparisons
@@
Afterward, we used the seven metrices to quantitively compare the results with other models, which are displayed in Table I. DAE-Fuse shows an outstanding performance across all the measurement indices, demonstrating the effectiveness of our method.
III-C Generalization on MIF tasks
To validate the generalizability of our DAE-Fuse, we used the same models in our IVIF testing to fuse medical images.
III-C1 Qualitative comparisons
Figure 2 compares the MRI-CT fusion results of different models. Similar to IVIF experiments, GAN-based models appear fused but are blurry and detail-poor. AE-based methods generally perform better in feature extraction. However, models like U2fusion and RFN-Nest underweight CT images, leading to darker outlines, whereas DIDFuse emphasizes CT features. Our method integrates both images effectively, preserving all texture details without bias, demonstrating superior generalization.
III-C2 Quantitative comparisons
Similarly, seven metrics are adopted to quantitatively compare the result, which are displayed in Table I. DAE-Fuse has the best score on almost all metrics, indicating that out method can be generalized to MIF tasks without any adjustment, and suitable for various kinds of MIF tasks.
III-D Downstream MMOD task
@@
Multi-Modality Object Detection is an important downstream task of image fusion. A single modality image i.e., an individual infrared image usually lacks certain features of objects during the detection process. As shown in Figure 4, infrared images exhibit a robust capability for detecting humans but may overlook objects that do not produce thermal radiation. On the other hand, visible images struggle to recognize humans due to reflective lights from vehicles and lamps. After fusing the images from two modalities, by combining the advantages of two types of features, both humans and vehicles are well detected in the fusion image.
Figure 4: Example of the distinct detection ability of the sourse images and our fused image.
IV Conclusion
In conclusion, DAE-Fuse overcomes limitations in image fusion, producing sharp and natural images through adversarial feature extraction and attention-guided fusion. Discriminative blocks in both phases enhance feature extraction and structural fidelity. Public dataset experiments demonstrate DAE-Fuse’s superiority over existing methods.
V Acknowledgement

### v2

@@ IV-C1 Qualitative comparisons
As shown in Figure 4, we select a representative scenario to evaluate feature extraction capabilities and highlight model biases. The GAN-based models DDcGAN and TarDAL exhibit a blurry style that loses many details, while the AE-based method DIDFuse shows a clear bias toward infrared images, rendering the sky and grass excessively dark. In contrast, DAE-Fuse achieves the best outcome, preserving rich texture details and maintaining a seamless balance between modalities. Notably, it fully retains the roof’s texture, a result unmatched by other methods. Although CDDFuse, with its parallel encoder architecture, maintains some roof details, parts are overexposed, and additional noise is introduced to the wall textures. By comparison, DAE-Fuse naturally fuses the walls from both inputs, demonstrating superior fusion quality.
IV-C2 Quantitative comparisons
@@
Afterward, we use the seven metrices to quantitively compare the results with other models, which are displayed in Table I. DAE-Fuse shows an outstanding performance across all the measurement indices, demonstrating the effectiveness of our method.
TABLE II: Quantitative comparisons on MRI-CT datasets. Bold red indicates the best, and bold blue indicates the second best.
Dataset: MRI-CT
EN SD SF MI SCD VIF Qabf DIDFuse 4.37 58.34 34.64 1.71 0.69 0.41 0.38 U2Fusion 4.21 61.98 32.54 2.08 0.75 0.37 0.46 RFN-Nest 4.97 70.36 33.42 1.98 0.68 0.43 0.52 DDcGAN 4.26 62.56 30.61 1.72 0.65 0.38 0.42 TarDAL 4.35 61.14 28.38 1.94 0.92 0.32 0.56 CDDFuse 4.49 71.36 34.02 2.16 1.18 0.44 0.56 DDFM 4.77 69.35 32.77 1.98 1.03 0.41 0.54 Ours 4.83 76.19 35.56 2.20 1.21 0.49 0.57
TABLE III: Quantitative comparison of our method with 7 state-of-the-art methods in AP@50(%) values for MM object detection on M3FD dataset. Bold red indicates the best, and Bold blue indicates the second best.
Peo Car Lam Bus Mot Tru mAP
Ir 0.804 0.886 0.712 0.802 0.725 0.743 0.779
Vis 0.721 0.865 0.848 0.811 0.794 0.791 0.805
DID 0.791 0.924 0.857 0.833 0.787 0.788 0.830
U2F 0.802 0.922 0.870 0.839 0.783 0.786 0.833
RFN 0.813 0.915 0.851 0.829 0.813 0.875 0.849
DDc 0.797 0.908 0.832 0.895 0.805 0.872 0.851
TarD 0.835 0.947 0.854 0.928 0.811 0.874 0.874
CDD 0.846 0.928 0.864 0.931 0.813 0.891 0.878
DDFM 0.837 0.926 0.869 0.927 0.809 0.882 0.875
Ours 0.855 0.931 0.874 0.949 0.822 0.890 0.887
IV-D Object Detection Ability
@@
A single modality image i.e., an individual infrared image usually lacks certain features of objects during the detection process. As shown in Figure 3., infrared images exhibit a robust capability for detecting humans but may overlook objects that do not produce thermal radiation. On the other hand, visible images struggle to recognize humans due to reflective lights from vehicles and lamps. After fusing the images from two modalities, by combining the advantages of two types of features, both humans and vehicles are well detected in the fusion image. Table III exhibits the detection results. Benefiting from the excellent fusion ability on the thermal radiation information, DAE-Fuse successfully detected all the targets and outperformed all other SOTA methods.
IV-E Generalization on Medical Image Fusion tasks
To validate the generalizability and robustness of our DAE-Fuse, we used the same models in our IVIF testing to fuse medical images.
IV-E1 Performance
The qualitative and quantitative comparisons for source images and fusion results on the MRI-CT dataset are shown in Figure 5 and Table II, respectively. Similar to the IVIF experiments, GAN-based models produce blurry results with significant detail loss, while AE-based methods generally perform better in feature extraction. For example, U2Fusion and RFN-Nest assign lower weight to CT images, leading to darker target outlines, while DIDFuse emphasizes CT features more conspicuously. In contrast, DAE-Fuse achieves superior integration of information from both modalities, preserving critical details and maintaining a balanced fusion. Quantitatively, DAE-Fuse outperforms other models on nearly all metrics, demonstrating its generalizability and adaptability to MIF tasks without requiring adjustments.
IV-F Ablation study
The ablation study is performed on the TNO dataset. We use EN, SD, VIF and Qabf to present the quantitative result, which is illustrated in Table LABEL:tab:ablation.
IV-F1 Discriminative blocks
In this experiment, we removed the discriminator blocks while keeping all other configurations unchanged. In phase one, performance decreased across all metrics, highlighting the role of discriminative blocks in guiding feature extraction and improving image quality. Similarly, in phase two, their contribution to the final outcomes demonstrates the importance of the adversarial game between the fused results and input pairs.
IV-F2 Cross-attention module
To verify the contribution of the cross-attention module, we conducted an experiment by abandoning it and using only the channel concatenation in phase two with the discriminative blocks reserved. The result shows that all the matrices decrease. The interaction between the embeddings of two modalities brought by the cross-attention did boost the fusion results.
TABLE IV: Ablation comparisons of our method on TNO dataset. Bold red indicates the best.
Configurations EN SD VIF Qabf
w/o discriminative blocks in phase one 6.81 43.20 0.91 0.63
w/o discriminative blocks in phase two 6.78 45.31 0.83 0.61
w/o cross-attention module 6.78 43.79 0.89 0.62
Ours 6.86 44.05 0.98 0.68
V Conclusion
In conclusion, our novel DAE-Fuse framework, tailored for autonomous navigation, effectively overcomes the limitations of traditional image fusion methods by generating sharp and natural fused images. This is accomplished through a two-phase approach: adversarial feature extraction and attention-guided cross-modality fusion. The integration of discriminative blocks and a cross-attention module within our encoder-decoder architecture ensures enhanced feature extraction and maintains structural integrity, resulting in visually appealing fused images. Evaluations on diverse datasets confirm DAE-Fuse’s superiority, not only in standard metrics but also in improving object detection for autonomous vehicles, setting a new standard in image fusion for real-world applications.
VI Acknowledgement